# All-in-One: Heterogeneous Interaction Modeling for Cold-Start Rating Prediction

### v2

@@ IV-A Overview
The gist of HIRE is to use a simple yet effective mechanism, MHSA, to holistically learn the interactions in three levels, i.e., users, items, and attributes, in an end-to-end, and data-driven fashion. The overall architecture of HIRE is illustrated in Fig. 3. To predict the rating of a user on an item, where either the user or item or both is cold-start, first we generate a prediction context which is composed of the target user/item, together with a set of pertinent users and items. These pertinent users and items are sampled from the user-item bipartite graph by leveraging the rating interaction. We envision that such users and items randomly selected by explicit rating relationships are informative for the cold-start user/item. Second, for a sampled prediction context, we construct a context matrix as the input of the model, which is a 3-dimensional tensor account for the users, items and attributes. The attributes contain individual attributes of users and items, and the observed ratings between the users and items in the context, where the ratings to be predicted are masked. The context matrix is transformed by $K$ Heterogeneous Interaction Module (HIM) blocks, the main component of the HIRE model, followed by a decoder to generate the prediction matrix. In each HIM block, there are three MHSA layers that learn the interactions between different users, different items and different categories of attributes, respectively. Intuitively, as Fig. 4 shows, the three MHSA layers learn via message passing in three complete graphs, where the nodes are the users, items in the context, and the involved attributes, respectively. In the rest of this section, we will elaborate on the details of the construction of the prediction context and the HIM block in § IV-B and § IV-C, respectively.
From the perspective of the prediction task, what a HIRE model does is analogous to inductive matrix completion [43, 44], where Graph Neural Networks are used to model the interactions in a matrix. However, from the perspective of model design, HIRE is different from these GNN-based approaches. For one thing, GNN-based approaches only learn from a single type of rating relationship. For the other thing, the MHSA we use is a flexible specification of GNN layer [45]. As Eq. (2)-(3) shows, self-attention conducts the message passing on a complete graph with a learned ’soft’ adjacency matrix instead of a fixed adjacency matrix.
@@ IV-B Construction of Prediction Context
$\bm{x}_{u}$ $\leftarrow[f_{U}^{1}(\bm{e}_{u}^{1})\|f_{U}^{2}(\bm{e}_{u}^{2})\|\dots\|f_{U}^{h_{u}}(\bm{e}_{u}^{h_{u}})]$ (7)
$\bm{x}_{i}$ $\leftarrow[f_{I}^{1}(\bm{e}_{i}^{1})\|f_{I}^{2}(\bm{e}_{i}^{2})\|\dots\|f_{I}^{h_{i}}(\bm{e}_{i}^{h_{i}})]$ (8)
For the ratings, we also use a linear transformation $f_{R}$ to map the concrete one-hot encoding of a rating $r$, $\bm{e}_{r}$, to the hidden representation $\bm{x}_{r}$ as shown in Eq. (9).
$\bm{x}_{r}\leftarrow f(\bm{e}_{r})$ (9)
@@ IV-C Model Details
Modeling interactions between users (MBU). Given the initial embedding of the context $\bm{H}\in\mathbb{R}^{n\times m\times e}$, where $n$ is the number of users, $m$ is the number of items, $e$ is the embedding dimension of attributes, we introduce an attention layer between users to model their implicit interactions. Specifically, we use $H_{i_{j}}=\bm{H}[\cdot,j,\cdot]\in\mathbb{R}^{n\times e}$ to denote the embedding view of item $j$ in $\bm{H}$. An item embedding view, $H_{i_{j}}$, is processed by a multi-head self-attention (MHSA) layer, where the attention weights between $n$ users are computed, which is specific to item $j$. In Eq. (10), $\tilde{H}_{i_{j}}\in\mathbb{R}^{m\times e}$ is the fused item embedding view weighted by the attention weights among users. All the item embedding views are processed by a parameter-sharing MHSA in parallel, and the fused embeddings, $\{\tilde{H}_{i_{j}}|j\in[1,m]\}$, are stacked as the output embedding $\bm{H}^{(U)}\in\mathbb{R}^{n\times m\times e}$ (Eq. (11)).
$\tilde{H}_{i_{j}}$ $\leftarrow\text{MHSA}(H_{i_{j}}),\forall{j}\in[1,\cdots,m]$ (10)
$\bm{H}^{(U)}$ $\leftarrow[\tilde{H}_{i_{1}}\|\tilde{H}_{i_{2}}\|\cdots\|\tilde{H}_{i_{m}}]$ (11)
@@
Modeling interaction between items (MBI). To further model the attention between items, we introduce an attention layer between items, given the stacked output embedding $\bm{H}^{(U)}\in\mathbb{R}^{n\times m\times e}$ as the input. Similarly, ${H}_{u_{k}}=\bm{H}^{(U)}[k,\cdot,\cdot]\in\mathbb{R}^{m\times e}$ denotes the embedding view of user $k$ in $\bm{H}^{(U)}$. The attention weights between $m$ items are computed through a MHSA layer, which is specific to a user embedding view, $H_{u_{k}}$, indicating the user-specific interactions of items. As Eq. (12) shows, $\tilde{H}_{u_{k}}$ is the fused user embedding view weighted by the attention weights among items. A parameter-sharing MHSA processes $m$ independent user embedding views in parallel. The fused embeddings, $\{\tilde{H}_{u_{k}}|k\in[1,n]\}$, are stacked as the output embedding $\bm{H}^{(I)}\in\mathbb{R}^{n\times m\times e}$ (Eq. (13)).
$\tilde{H}_{u_{k}}$ $\leftarrow\text{MHSA}(H_{u_{k}}),\forall{k}\in[1,\cdots,n]$ (12)
$\bm{H}^{(I)}$ $\leftarrow[\tilde{H}_{u_{1}}\|\tilde{H}_{u_{2}}\|\cdots\|\tilde{H}_{u_{n}}]$ (13)
@@
Modeling interaction between attributes (MBA). HIM further learns the interaction between fine-grained attributes by computing the attention between user and item attributes. Taking the embedding $\bm{H}^{(I)}\in\mathbb{R}^{n\times m\times e}$ as input, we first reshape $\bm{H}^{(I)}$ to $\mathbb{R}^{n\times m\times h\times f}$ by spliting the feature dimension $e$ into $h\times f$, where $h$ is the number of categorical attributes and $f$ is the embedding dimension for each attribute. We use ${H}_{u_{k},i_{j}}=\bm{H}^{(I)}[k,j,\cdot,\cdot]\in\mathbb{R}^{h\times f}$ to denote the embedding view of a user-item pair $(u_{k},i_{j})$ in $\bm{H}^{(I)}$. The embedding view $H_{u_{k},i_{j}}$ is processed by a MHSA layer, where the attention weights between $h$ attributes are computed, indicating the interaction between attributes that is specific to a user-item pair $(u_{k},i_{j})$. $\tilde{H}_{u_{k},i_{j}}\in\mathbb{R}^{h\times f}$ in Eq. (14) is the fused user-item pair embedding view weighted by the attention weights between the $h$ attributes. All the user-item pair embedding views are processed by a parameter-sharing MHSA in parallel, and the fused embeddings, $\{\tilde{H}_{u_{k},i_{j}}|k\in[1,n],j\in[1,m]\}$, are stacked as the output embedding $\bm{H}^{(A)}\in\mathbb{R}^{n\times m\times h\times f}$ (Eq. (15)). The embedding $\bm{H}^{(A)}$ will be reshaped to $\mathbb{R}^{n\times m\times e}$ by flatting the feature dimension $h\times f$ back to $e$ for future processing.
$\tilde{H}_{u_{k},i_{j}}$ $\leftarrow\text{MHSA}(H_{u_{k},i_{j}}),\forall{k}\in[1,\cdots,n],\forall{j}\in[1,\cdots,m]$ (14)
@@
$\tilde{H}_{u_{n},i_{1}}\|\tilde{H}_{u_{n},i_{2}}\|\cdots\|\tilde{H}_{u_{n},i_{m}}]$ (15)
To summarize, HIM takes the sampled context $\bm{H}\in\mathbb{R}^{n\times m\times e}$ as input embedding, and model the interactions between users, items and attributes by three MHSA layer by layer. It finally generates embedding $\bm{H}^{(A)}\in\mathbb{R}^{n\times m\times e}$ which adaptively aggregates the features of users, items and attributes in the context by leveraging the attention weights. As Fig. 3 presents, the model HIRE is composed of $K$ HIMs, where the output of the $(k-1)$-th HIM is fed into $k$-th HIM for modeling the higher order interactions.
Rating Prediction. Given $\bm{H}^{(A)}\in\mathbb{R}^{n\times m\times e}$ as the output of the $K$-th HIM, we use a decoder to predict the rating matrix $\hat{R}\in\mathbb{R}^{n\times m}$ of the sampled context as Eq. (16).
$\hat{R}\leftarrow\alpha\cdot\text{Sigmoid}(g_{\theta}(\bm{H}^{(A)}))$ (16)
@@ V-A Model Training
In a recommendation system, given a set of users ${\mathcal{U}}$, a set of items ${\mathcal{I}}$ and observed ratings ${\mathcal{R}}$, a HIRE model is trained by a set of prediction contexts, denoted as $\mathcal{D}$, which are sampled from ${\mathcal{U}}$, ${\mathcal{I}}$, and ${\mathcal{R}}$. Given a ratio of $p$ ratings in a prediction context, the model is optimized to minimize the mean squared error (MSE) loss (Eq. (17)) for the $1-p$ of masked ratings.
$\mathcal{L}=\frac{1}{|\mathcal{D}|}\sum_{\mathcal{D}}\frac{1}{|\mathcal{Q}|}\sum_{r\in\mathcal{Q}}\norm{r-\hat{r}}^{2}$ (17)
@@
Algorithm 1 presents the training process implemented by stochastic gradient descent. In each step, we randomly draw a mini-batch of prediction context, $\mathcal{B}$, from the training set $\mathcal{D}$ as line 1. Then, in line 1-1, each context is transformed by the three attention layers in a HIM block one by one, and there are $K$ HIM blocks in total. In line 1, the output of the $K$-th HIM is mapped to the predicted rating matrix $\hat{R}$. Finally, we compute the MSE Loss for the mini-batch $\mathcal{B}$ and update the model parameters via the back-propagation of gradient in line 1-1. The training will be terminated until the MSE loss converges.
In test stage, when new user set $\mathcal{U}^{*}$ and item set $\mathcal{I}^{*}$ comes, we can construct an embedding matrix $\bm{H}^{*}$. Ratings to be predicted are masked. It is worth to mention that the size of matrix can be decided by the number of new users and items and can be flexible. $\bm{H}^{*}$ is input into $K$ HIM blocks following a mapping function to get predicted rating.
@@ V-B Model Analysis
Property 5.1. Given a set of users $\{u_{1},\cdots,u_{n}\}$ and a set of items $\{i_{1},\cdots,i_{m}\}$ as a prediction context $\bm{H}$ and the HIRE model $\mathcal{M}$, the predicted rating matrix, $\hat{R}$, is equivariant w.r.t. any permutation of the user set $\Pi_{[1:n]}$ and permutation of the item set $\Pi_{[1:m]}$.
$\Pi_{[1:n]}\circ\Pi_{[1:m]}\circ\hat{R}=\mathcal{M}(\Pi_{[1:n]}\circ\Pi_{[1:m]}\circ\bm{H})$ (18)
@@
$\Pi_{[1:n]}\circ\tilde{H}_{i_{j}}$ $\leftarrow\text{MHSA}(\Pi_{[1:n]}\circ H_{i_{j}})$ (19)
$\Pi_{[1:m]}\circ\tilde{H}_{u_{k}}$ $\leftarrow\text{MHSA}(\Pi_{[1:m]}\circ H_{u_{k}})$ (20)
@@ 17       Update model parameters by gradient descent;
18
19while $\mathcal{L}$ converges
@@
TABLE II: Profile of Datasets
@@
TABLE III: Overall Performance in Three Cold-Start Scenarios on $\mathsf{MovieLens}$-$\mathsf{1M}$

### v3

@@ IV-A Overview
The gist of HIRE is to use a simple yet effective mechanism, MHSA, to holistically learn the interactions in three levels, i.e., users, items, and attributes, in an end-to-end, and data-driven fashion. The overall architecture of HIRE is illustrated in Fig. 3. To predict the rating of a user on an item, where either the user or item or both is cold-start, first we generate a prediction context which is composed of the target user/item, together with a set of pertinent users and items. These pertinent users and items are sampled from the user-item bipartite graph by leveraging the rating interaction. We envision that such users and items randomly selected by explicit rating relationships are informative for the cold-start user/item. Second, for a sampled prediction context, we construct a context matrix as the input of the model, which is a 3-dimensional tensor account for the users, items and attributes. The attributes contain individual attributes of users and items, and the observed ratings between the users and items in the context, where the ratings to be predicted are masked. The context matrix is transformed by $L$ Heterogeneous Interaction Module (HIM) blocks, the main component of the HIRE model, followed by a decoder to generate the prediction matrix. In each HIM block, there are three MHSA layers that learn the interactions between different users, different items and different categories of attributes, respectively. Intuitively, as Fig. 4 shows, the three MHSA layers learn via message passing in three complete graphs, where the nodes are the users, items in the context, and the involved attributes, respectively. In the rest of this section, we will elaborate on the construction of the prediction context and the HIM in § IV-B and § IV-C, respectively.
From the perspective of the prediction task, what a HIRE model does is analogous to inductive matrix completion [52, 53], where Graph Neural Networks are used to model the interactions in a matrix. However, from the perspective of model design, HIRE is different from these GNN-based approaches. For one thing, GNN-based approaches only learn from a single type of rating relationship. For the other thing, the MHSA we use is a flexible specification of GNN layer [54]. As Eq. (2)-(3) shows, self-attention conducts the message passing on a complete graph with a learned ‘soft’ adjacency matrix instead of a fixed adjacency matrix.
@@ IV-B Construction of Prediction Context
$\bm{x}_{u}$ $\leftarrow[f_{U}^{1}(\bm{e}_{u}^{1})\|f_{U}^{2}(\bm{e}_{u}^{2})\|\dots\|f_{U}^{h_{u}}(\bm{e}_{u}^{h_{u}})],$ (7)
$\bm{x}_{i}$ $\leftarrow[f_{I}^{1}(\bm{e}_{i}^{1})\|f_{I}^{2}(\bm{e}_{i}^{2})\|\dots\|f_{I}^{h_{i}}(\bm{e}_{i}^{h_{i}})].$ (8)
For the ratings, we also use a linear transformation $f_{R}$ to map the concrete one-hot encoding of a rating $r$, $\bm{e}_{r}$, to the hidden representation $\bm{x}_{r}$ as shown in Eq. (9),
$\bm{x}_{r}\leftarrow f(\bm{e}_{r}),$ (9)
@@ IV-C Model Details
Modeling interactions between users (MBU). Given the initial embedding of the context $\bm{H}\in\mathbb{R}^{n\times m\times e}$, where $n$ is the number of users, $m$ is the number of items, $e$ is the embedding dimension of attributes, we introduce an attention layer between users to model their implicit interactions. Specifically, we use $H_{i_{j}}=\bm{H}[\cdot,j,\cdot]\in\mathbb{R}^{n\times e}$ to denote the embedding view of item $j$ in $\bm{H}$. An item embedding view, $H_{i_{j}}$, is processed by a multi-head self-attention (MHSA) layer, where the attention weights between $n$ users are computed, which is specific to item $j$. In Eq. (10), $\tilde{H}_{i_{j}}\in\mathbb{R}^{m\times e}$ is the fused item embedding view weighted by the attention weights among users. All the item embedding views are processed by a parameter-sharing MHSA in parallel, and the fused embeddings, $\{\tilde{H}_{i_{j}}|j\in[1,m]\}$, are stacked as the output embedding $\bm{H}^{(U)}\in\mathbb{R}^{n\times m\times e}$ (Eq. (11)),
$\tilde{H}_{i_{j}}$ $\leftarrow\text{MHSA}(H_{i_{j}}),\forall{j}\in[1,\cdots,m],$ (10)
$\bm{H}^{(U)}$ $\leftarrow[\tilde{H}_{i_{1}}\|\tilde{H}_{i_{2}}\|\cdots\|\tilde{H}_{i_{m}}].$ (11)
@@
Modeling interaction between items (MBI). To further model the interaction between items, we introduce an attention layer between items, given the stacked output embedding $\bm{H}^{(U)}\in\mathbb{R}^{n\times m\times e}$ as the input. Similarly, ${H}_{u_{k}}=\bm{H}^{(U)}[k,\cdot,\cdot]\in\mathbb{R}^{m\times e}$ denotes the embedding view of user $k$ in $\bm{H}^{(U)}$. The attention weights between $m$ items are computed through a MHSA layer, which is specific to a user embedding view, $H_{u_{k}}$, indicating the user-specific interactions of items. As Eq. (12) shows, $\tilde{H}_{u_{k}}$ is the fused user embedding view weighted by the attention weights among items. A parameter-sharing MHSA processes $m$ independent user embedding views in parallel. The fused embeddings, $\{\tilde{H}_{u_{k}}|k\in[1,n]\}$, are stacked as the output embedding $\bm{H}^{(I)}\in\mathbb{R}^{n\times m\times e}$ (Eq. (13)),
$\tilde{H}_{u_{k}}$ $\leftarrow\text{MHSA}(H_{u_{k}}),\forall{k}\in[1,\cdots,n],$ (12)
$\bm{H}^{(I)}$ $\leftarrow[\tilde{H}_{u_{1}}\|\tilde{H}_{u_{2}}\|\cdots\|\tilde{H}_{u_{n}}].$ (13)
@@
Modeling interaction between attributes (MBA). HIM further learns the interaction between fine-grained attributes by computing the attention between user and item attributes. Taking the embedding $\bm{H}^{(I)}\in\mathbb{R}^{n\times m\times e}$ as input, we first reshape $\bm{H}^{(I)}$ to $\mathbb{R}^{n\times m\times h\times f}$ by spliting the feature dimension $e$ into $h\times f$, where $h$ is the number of categorical attributes and $f$ is the embedding dimension for each attribute. We use ${H}_{u_{k},i_{j}}=\bm{H}^{(I)}[k,j,\cdot,\cdot]\in\mathbb{R}^{h\times f}$ to denote the embedding view of a user-item pair $(u_{k},i_{j})$ in $\bm{H}^{(I)}$. The embedding view $H_{u_{k},i_{j}}$ is processed by a MHSA layer, where the attention weights between $h$ attributes are computed, indicating the interaction between attributes that is specific to a user-item pair $(u_{k},i_{j})$. $\tilde{H}_{u_{k},i_{j}}\in\mathbb{R}^{h\times f}$ in Eq. (14) is the fused user-item pair embedding view weighted by the attention weights between the $h$ attributes. All the user-item pair embedding views are processed by a parameter-sharing MHSA in parallel, and the fused embeddings, $\{\tilde{H}_{u_{k},i_{j}}|k\in[1,n],j\in[1,m]\}$, are stacked as the output embedding $\bm{H}^{(A)}\in\mathbb{R}^{n\times m\times h\times f}$ (Eq. (15)). The embedding $\bm{H}^{(A)}$ will be reshaped to $\mathbb{R}^{n\times m\times e}$ by flatting the feature dimension $h\times f$ back to $e$ for future processing,
$\tilde{H}_{u_{k},i_{j}}$ $\leftarrow\text{MHSA}(H_{u_{k},i_{j}}),\forall{k}\in[1,\cdots,n],\forall{j}\in[1,\cdots,m],$ (14)
@@
$\tilde{H}_{u_{n},i_{1}}\|\tilde{H}_{u_{n},i_{2}}\|\cdots\|\tilde{H}_{u_{n},i_{m}}].$ (15)
To summarize, HIM takes the sampled context $\bm{H}\in\mathbb{R}^{n\times m\times e}$ as input embedding, and model the interactions between users, items and attributes by three MHSA layer by layer. It finally generates embedding $\bm{H}^{(A)}\in\mathbb{R}^{n\times m\times e}$ which adaptively aggregates the features of users, items and attributes in the context by leveraging the attention weights. As Fig. 3 presents, the model HIRE is composed of $L$ HIMs, where the output of the $(l-1)$-th HIM is fed into $l$-th HIM for modeling the higher order interactions.
Rating Prediction. Given $\bm{H}^{(A)}\in\mathbb{R}^{n\times m\times e}$ as the output of the $L$-th HIM, we use a decoder to predict the rating matrix $\hat{R}\in\mathbb{R}^{n\times m}$ of the sampled context as Eq. (16),
$\hat{R}\leftarrow\alpha\cdot\text{Sigmoid}(g_{\theta}(\bm{H}^{(A)})).$ (16)
@@ V-A Model Training
In a recommendation system, given a set of users ${\mathcal{U}}$, a set of items ${\mathcal{I}}$ and observed ratings ${\mathcal{R}}$, a HIRE model is trained by a set of prediction contexts, denoted as $\mathcal{D}$, which are sampled from ${\mathcal{U}}$, ${\mathcal{I}}$, and ${\mathcal{R}}$. Given a ratio of $p$ ratings in a prediction context, the model is optimized to minimize the mean squared error (MSE) loss (Eq. (17)) for the $1-p$ of masked ratings,
$\mathcal{L}=\frac{1}{|\mathcal{D}|}\sum_{\mathcal{D}}\frac{1}{|\mathcal{Q}|}\sum_{r\in\mathcal{Q}}\norm{r-\hat{r}}^{2},$ (17)
@@
Algorithm 1 presents the training process implemented by stochastic gradient descent. In each step, we randomly draw a mini-batch of prediction context, $\mathcal{B}$, from the training set $\mathcal{D}$ as line 1. Subsequently, in line 1-1, each context is transformed by the three attention layers in a HIM block one by one, and there are $L$ HIM blocks in total. In line 1, the output of the $L$-th HIM is mapped to the predicted rating matrix $\hat{R}$. Finally, we compute the MSE Loss for the mini-batch $\mathcal{B}$ and update the model parameters via the back-propagation of gradient in line 1-1. The training will be terminated until the MSE loss converges.
In the test stage, when new user set $\mathcal{U}^{*}$ and item set $\mathcal{I}^{*}$ arrives, we can construct an embedding matrix $\bm{H}^{*}$. Ratings to be predicted are masked as 00. $\bm{H}^{*}$ is input into $L$ HIM blocks following a mapping function to get predicted rating.
Example 2
Given the prediction context $\{u_{1},u_{2},i_{1},i_{2}\}$ in Example 1, we suppose both users and items have 2 attributes. The context matrix $\bm{H}$ is a ${2\times 2\times 80}$ tensor, with each attribute and the rating are transformed into 16-dimensional vector by Eq. (7)-(9). Specifically, the embedded ratings contain the observed rating of $(u_{2},i_{1})$ and $(u_{2},i_{2})$, the target rating of $(u_{1},i_{2})$ which is masked as 00, and the unobserved rating of $(u_{1},i_{1})$ which is masked as $-1$. Taking $\bm{H}$ as the model input, MBU, MBI, MBA compute the embeddings $\bm{H}^{(U)},\bm{H}^{(I)},\bm{H}^{(A)}\in\mathbb{R}^{2\times 2\times 80}$, respectively. The process is repeated by $L$ times. Finally, the last $\bm{H}^{(A)}$ is fed into a decoder in Eq. (16) to generate the rating matrix $\hat{R}\in\mathbb{R}^{2\times 2}$. The predicted rating of $(u_{1},i_{2})$ corresponds to the element in the first row and second column of the matrix, i.e., $\hat{R}[0,1]$.
@@ V-B Model Analysis
Property 5.1. Given a set of users $\{u_{1},\cdots,u_{n}\}$ and a set of items $\{i_{1},\cdots,i_{m}\}$ as a prediction context $\bm{H}$ and the HIRE model $\mathcal{M}$, the predicted rating matrix, $\hat{R}$, is equivariant w.r.t. any permutation of the user set $\Pi_{[1:n]}$ and permutation of the item set $\Pi_{[1:m]}$,
$\Pi_{[1:n]}\circ\Pi_{[1:m]}\circ\hat{R}=\mathcal{M}(\Pi_{[1:n]}\circ\Pi_{[1:m]}\circ\bm{H}).$ (18)
@@
$\Pi_{[1:n]}\circ\tilde{H}_{i_{j}}$ $\leftarrow\text{MHSA}(\Pi_{[1:n]}\circ H_{i_{j}}),$ (19)
$\Pi_{[1:m]}\circ\tilde{H}_{u_{k}}$ $\leftarrow\text{MHSA}(\Pi_{[1:m]}\circ H_{u_{k}}),$ (20)
@@ 15       Update model parameters by gradient descent;
16
17while $\mathcal{L}$ converges
@@
TABLE I: Profile of Datasets
@@
TABLE II: Overall Performance in Warm States (%)
Datasets Methods TOP5 TOP7 TOP10
Precision NDCG MAP Precision NDCG MAP Precision NDCG MAP
$\mathsf{MovieLens}$-$\mathsf{1M}$ Popularity $57.63_{\small(0.52)}$ $86.02_{\small(0.13)}$ $47.41_{\small(0.21)}$ $62.47_{\small(0.63)}$ $86.64_{\small(0.11)}$ $51.74_{\small(0.38)}$ $69.69_{\small(0.63)}$ $87.80_{\small(0.08)}$ $59.63_{\small(0.58)}$
NeuMF $49.92_{\small(0.12)}$ $73.01_{\small(0.11)}$ $40.83_{\small(0.10)}$ $52.25_{\small(0.19)}$ $74.97_{\small(0.15)}$ $42.26_{\small(0.16)}$ $56.01_{\small(0.20)}$ $76.62_{\small(0.17)}$ $45.71_{\small(0.21)}$
TaNP $58.73_{\small(0.12)}$ $87.24_{\small(0.06)}$ $48.66_{\small(0.04)}$ $61.79_{\small(0.09)}$ $87.55_{\small(0.05)}$ $50.95_{\small(0.20)}$ $67.21_{\small(0.17)}$ $88.27_{\small(0.02)}$ $56.56_{\small(0.24)}$
HIRE $\mathbf{78.23_{\small(0.57)}}$ $\mathbf{95.86_{\small(0.26)}}$ $\mathbf{73.06_{\small(0.68)}}$ $\mathbf{78.46_{\small(0.37)}}$ $\mathbf{96.58_{\small(0.24)}}$ $\mathbf{73.60_{\small(0.46)}}$ $\mathbf{78.51_{\small(0.13)}}$ $\mathbf{98.03_{\small(0.22)}}$ $\mathbf{73.90_{\small(0.16)}}$
$\mathsf{Bookcrossing}$ Popularity $35.92_{\small(0.19)}$ $83.02_{\small(0.12)}$ $25.58_{\small(0.23)}$ $41.05_{\small(0.28)}$ $83.83_{\small(0.10)}$ $27.75_{\small(0.31)}$ $50.05_{\small(0.28)}$ $85.47_{\small(0.07)}$ $34.36_{\small(0.28)}$
NeuMF $40.69_{\small(0.13)}$ $33.07_{\small(0.16)}$ $36.49_{\small(0.68)}$ $53.02_{\small(0.12)}$ $34.34_{\small(0.19)}$ $48.83_{\small(0.09)}$ $65.96_{\small(0.07)}$ $37.33_{\small(0.17)}$ $62.01_{\small(0.05)}$
TaNP $51.47_{\small(0.55)}$ $88.33_{\small(0.19)}$ $44.11_{\small(0.59)}$ $55.15_{\small(0.54)}$ $88.89_{\small(0.16)}$ $46.53_{\small(0.50)}$ $61.84_{\small(0.27)}$ $89.89_{\small(0.12)}$ $52.02_{\small(0.60)}$
HIRE $\mathbf{60.86_{\small(1.22)}}$ $\mathbf{91.09_{\small(0.44)}}$ $\mathbf{55.00_{\small(1.53)}}$ $\mathbf{65.51_{\small(0.70)}}$ $\mathbf{92.42_{\small(0.44)}}$ $\mathbf{58.54_{\small(0.82)}}$ $\mathbf{70.40_{\small(0.21)}}$ $\mathbf{94.47_{\small(0.42)}}$ $\mathbf{63.85_{\small(0.45)}}$
$\mathsf{Douban}$ Popularity $57.28_{\small(0.18)}$ $87.61_{\small(0.01)}$ $47.54_{\small(0.13)}$ $63.74_{\small(0.22)}$ $88.51_{\small(0.03)}$ $53.50_{\small(0.25)}$ $73.66_{\small(0.03)}$ $90.14_{\small(0.01)}$ $64.50_{\small(0.06)}$
NeuMF $35.45_{\small(0.05)}$ $43.32_{\small(0.04)}$ $28.29_{\small(0.02)}$ $41.90_{\small(0.05)}$ $44.30_{\small(0.04)}$ $34.44_{\small(0.02)}$ $53.81_{\small(0.05)}$ $44.75_{\small(0.05)}$ $47.23_{\small(0.04)}$
TaNP $57.26_{\small(1.02)}$ $86.63_{\small(0.22)}$ $46.34_{\small(0.82)}$ $66.15_{\small(0.94)}$ $88.17_{\small(0.24)}$ $55.10_{\small(0.91)}$ $\mathbf{78.94_{\small(1.10)}}$ $90.91_{\small(0.30)}$ $70.81_{\small(1.37)}$
HIRE $\mathbf{64.05_{\small(0.41)}}$ $\mathbf{90.07_{\small(0.23)}}$ $\mathbf{58.16_{\small(0.57)}}$ $\mathbf{69.39_{\small(0.35)}}$ $\mathbf{91.33_{\small(0.21)}}$ $\mathbf{62.73_{\small(0.55)}}$ ${76.70_{\small(0.20)}}$ $\mathbf{93.88_{\small(0.18)}}$ $\mathbf{71.10_{\small(0.35)}}$
TABLE III: Overall Performance in 3 Cold-Start Scenarios on $\mathsf{MovieLens}$-$\mathsf{1M}$ (%)